# Light-ResKAN: A Parameter-Sharing Lightweight KAN with Gram Polynomials for Efficient SAR Image Recognition

### 02 Apr 2026 v1

@@ V-E Few-Shot Recognition Experiments
AlexNet [alexnet] 38.48 46.95 57.64
VGG16 [vgg] 37.26 45.36 55.26
ResNet18 [resnet] 41.57 52.57 59.51
ResNet50 [resnet] 38.58 51.26 58.53
ConvNeXt [convnext] 27.36 38.02 51.37
MACN [macn] 21.94 30.93 45.33
Cross-HL [cross-hl] 38.23 46.52 57.31
ASANet [asanet] 24.74 33.96 47.58
@@ VI Conclusions and Future Work
Future research will focus on three strategic directions to advance Light-ResKAN. First, to bridge the gap between theoretical efficiency and deployment speed, we will implement optimization strategies such as CUDA kernel fusion and pursue deeper integration with mainstream deep learning frameworks. Second, to enhance representational power and address few-shot limitations, future work will investigate: (1) few-shot-aware polynomial regularization that adaptively constrains polynomial complexity based on training data volume; (2) integration with meta-learning frameworks (e.g., Prototypical Networks, MAML) for rapid polynomial coefficient adaptation; (3) self-supervised pre-training on unlabeled SAR imagery to improve polynomial parameter initialization. Furthermore, while the current noise injection experiments provide initial evidence of robustness under interference conditions, we plan to expand robustness evaluations by incorporating the SAR interference dataset [Yang2025SAR] to address real-world strong-interference scenarios—such as electronic countermeasures and severe clutter—while leveraging the large-scale ATRNet-STAR benchmark [liu2025atrnet] to assess generalization in realistic, complex environments. Finally, we aim to achieve deeper integration with mainstream deep learning frameworks to facilitate wider adoption and unlock potential for improved performance across a broader spectrum of SAR applications.

### 03 Apr 2026 v2

@@ V-E Few-Shot Recognition Experiments
AlexNet [19] 38.48 46.95 57.64
VGG16 [44] 37.26 45.36 55.26
ResNet18 [13] 41.57 52.57 59.51
ResNet50 [13] 38.58 51.26 58.53
ConvNeXt [34] 27.36 38.02 51.37
MACN [55] 21.94 30.93 45.33
Cross-HL [41] 38.23 46.52 57.31
ASANet [66] 24.74 33.96 47.58
@@ VI Conclusions and Future Work
Future research will focus on three strategic directions to advance Light-ResKAN. First, to bridge the gap between theoretical efficiency and deployment speed, we will implement optimization strategies such as CUDA kernel fusion and pursue deeper integration with mainstream deep learning frameworks. Second, to enhance representational power and address few-shot limitations, future work will investigate: (1) few-shot-aware polynomial regularization that adaptively constrains polynomial complexity based on training data volume; (2) integration with meta-learning frameworks (e.g., Prototypical Networks, MAML) for rapid polynomial coefficient adaptation; (3) self-supervised pre-training on unlabeled SAR imagery to improve polynomial parameter initialization. Furthermore, while the current noise injection experiments provide initial evidence of robustness under interference conditions, we plan to expand robustness evaluations by incorporating the SAR interference dataset [61] to address real-world strong-interference scenarios—such as electronic countermeasures and severe clutter—while leveraging the large-scale ATRNet-STAR benchmark [31] to assess generalization in realistic, complex environments. Finally, we aim to achieve deeper integration with mainstream deep learning frameworks to facilitate wider adoption and unlock potential for improved performance across a broader spectrum of SAR applications.